Find the value of the cosecant of an angle given that the point on the terminal side of an angle is (3,4)


1
Expert's answer
2022-02-22T00:45:52-0500

Solution:

Using the point (3,4), we can see that this forms a right triangle that has a base that is 3 units in length and an adjoining leg that is 4 units high. We are able to find the hypotenuse of this triangle using the Pythagorean Theorem. This will give us the distance of the point (3,4) to the origin.

So, "a=3,b=4"

Using,

"a^2+b^2=c^2
\\ \Rightarrow 3^2+4^2=c^2
\\\Rightarrow c^2=9+16=25
\\\Rightarrow c=5"

And so the hypotenuse of this triangle (the distance from our point we are working with to the origin), is 5 units long. Recall that when using cosecant for right triangles, cosecant represents the following

"\cosec=\dfrac{Hypotenuse}{Perpendicular}=\dfrac cb
\\\Rightarrow \cosec=\dfrac 54"